OpenStudy (anonymous):

a two digit number is obtained by either multiplying the sum of the digits by 8 and then subtracting 5 or by multiplying the difference of the digits by 16 and then adding 3.Find the number.

Parth (parthkohli):

It looks like you have a system to solve eh

Parth (parthkohli):

Let one digit be x and another one by y.

Parth (parthkohli):

\(\Rightarrow 8(x + y) - 5 = 10x + y \) \(\Rightarrow 16(x - y) + 3 = 10x + y \) Can you solve this one?

Parth (parthkohli):

Maybe, or you can try wolfram ;)

Parth (parthkohli):

I am not sure... I'd be going in some time

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

cn u solve it

OpenStudy (apoorvk):

@Parvathysubhash 9th grade - you should be able to do this stuff -_- Try once atleast :) Combine the like terms and you will have two linear equations to solve and enjoy! ;)

OpenStudy (anonymous):

dt i knww yaarr cn u jst show it

OpenStudy (anonymous):

8(x+y)-5=10x+y ----------------- (1) 16(x-y)+3=10x+y ___________________(2) (1) gives 8x+8y-5=10x+y (2)gives 16x-16y+3=10x+y after dt wht will we do @apoorvk

OpenStudy (apoorvk):

okay bring over the terms on the RHS to the left and simplify the whole thing.

OpenStudy (anonymous):

(1)- -2x-7y-5 (2)- 6x+17y+3

OpenStudy (anonymous):

nthng i dnt need d answer ok

OpenStudy (apoorvk):

there's a slight mistake: should read: -2x+7y-5 = 0 6x-17y+3 = 0 Now solve these two^